OpenStudy (anonymous):

What is the slope of the line that passes through the points (2,5) and (7,3)? A. -5/2 B. -2/5 C. 8/9 D. 9/8

OpenStudy (anonymous):

This is the formula for finding the slope of a line. m is the variable used for the slope. Slope = (y2-y1)/x2-x1 (2,5)and (7,3)? (2, 5) and (7,3)? (x1, y1) and (x2, y2) x1 = 2 and x2 = 7, y1 =5 and y2 = 3 Slope = (y2-y1)/x2-x1 = (3-5)/ (7-2) Now we are going to have: -2/5 Ans. -2 over 5

OpenStudy (anonymous):

Thank you Zoe! :)

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

glad to help :)